Calculate the pH and pOH for each of the solutions with the following hydrogen ion or hydroxide ion concentrations.

Solution B: [OH] = 2.56×10-4  M
Answer #1

B)
Given:
[OH-] = 2.56*10^-4 M

use:
pOH = -log [OH-]
= -log (2.56*10^-4)
= 3.592

Given:
pOH = 3.592

use:
PH = 14 - pOH
= 14 - 3.592
= 10.408

Answer:
pH = 10.408
pOH = 3.592